One number is 62 more than another number. If twice the greater number is subtracted from five times the lesser number, the result is 155. What are the two numbers?
Answer by mananth(16946)

x-y=62
5y-2x=155
1 x -1 y = 62 .............1
-2 x + 5 y = 155 .............2
Eliminate y
multiply (1)by 5
Multiply (2) by 1
5 x -5 y = 310
-2 x 5 y = 155
Add the two equations
3 x = 465
/ 3
x = 155
plug value of x in (1)
1 x + -1 y = 62
155 + -1 y = 62
-1 y = 62 -155
-1 y = -93
y = 93.00
155 ,93
m.ananth@hotmail.ca